can you name the property illustrated by each
1.x + y = y + x
2.k + 0 = k

Answer by stanbon(75887)

can you name the property illustrated by each
1.x + y = y + x
The communtative property of addition.
---------------------
2.k + 0 = k
The additive identity property.
-----------------------
Cheers,
Stan H.